LASER SPECTROSCOPY OF IRIDIUM MONOBORIDE

JIANJUN YE, H.F. PANG, A. M-Y. WONG, J. W-H. LEUNG AND \underline{A. S-C. CHEUNG}, Department of Chemistry, The University of Hong Kong, Pokfulam Road, Hong Kong.

High resolution laser induced fluorescence spectrum of IrB in the spectral region between 545 to 610 nm has been recorded and analyzed. Reacting laser-ablated iridium atoms with 1% B_{2}H_{6} seeded in argon produced the IrB molecule. Four vibronic transition bands: (v, 0) with v = 0-3 of an electronic transition system have been observed. Spectra of all four isotopic molecules: ^{191}Ir^{10}B, ^{193}Ir^{10}B, ^{191}Ir^{11}B and ^{193}Ir^{11}B were recorded. Relationship between isotopic molecules confirmed the vibrational quantum number assignment. Preliminary analysis of the rotational lines showed that these vibronic bands are with Ω′ = 2 and Ω′′ = 3. Partially resolved hyperfine structure conformed to case a_{β} coupling scheme has been observed.